OpenStudy (anonymous):

The attendance on the first day of a carnival was 425 people. The attendance on the third day was 575 people. Assuming attendance will increase linearly each day, how many people will attend the carnival on the sixth day? a)650 b)725 C)800 d)875

OpenStudy (anonymous):

suppose x=day y=total attendances on the first day (x=1) ,y=425 on the third day (x=2), y=575 the equation of a line is y=mx+b plug the value of x and y 425 = m + b 575 = 3m + b solve that , you'll get m=75 and b= 350 so the equation of the line is y = 75x + 350 if you want to find the number of attendances on sixth day, plug x=6 , you'll get the value of y(attendances) y = 75(6) + 350 = 800

OpenStudy (anonymous):

thanks love, your AMAZING!

OpenStudy (anonymous):

youre welcome